# Key varieties for prime $\mathbb{Q}$-Fano threefolds defined by Freudenthal triple systems

### v2

@@ 2.5. More on FTS and the scheme $\mathfrak{R}$
Explicit descriptions of the quartic form $\widetilde{F}$ and the symplectic form $\omega$ are given as follows (cf.[Cl, (21) in Sec.7, p.118–119],[Fe, p.321], [LM, Prop.5.5]). The proof is same as that given as in [Fe, p.321] with the above formulation.
@@
(2.27) $\Delta_{x}(s^{-1}y^{\sharp},y,\widetilde{x})=\Delta_{y}(s^{-1}y^{\sharp},y,\widetilde{y})=0\,\,\text{for any }\widetilde{p}=(\widetilde{s},\widetilde{t,}\widetilde{x},\widetilde{y}).$
@@ 2.6. A coordinatization of $\mathfrak{R}$ with parameters
Assume that $N_{x}$ is not identically zero on $V_{x}$. The equations (2.23) as in Theorem 2.18 is reduced to the two equations
@@
where $P,Q$ are $3\times 3$ $k$-matrices, and $x,y$ are considered as column vectors (${}^{t}\!(x_{1},x_{2},x_{3})$ and (${}^{t}\!(y_{1},y_{2},y_{3})$. Moreover, there exist nonzero constants $\mu,\nu\in\mathsf{k}$ such that
@@
and, by replacing $s,t$ with $\mu s,\nu$t respectively, the equation of $\mathfrak{R}$ is reduced to
@@
    There exist the uniquely determined cubics forms cubic forms $N_{x}^{PQ}(x)$ and $N_{y}^{PQ}(y)$ such that
@@
3. Affine scheme $\mathfrak{F}_{\mathbb{A}}^{22}$
@@
In the affine $26$-space whose coordinates are $s,t,$ and the entries of $x,y,P,Q$, we define $\mathfrak{F}_{\mathbb{A}}^{22}$ to be the scheme with the equations (2.33), (2.34) and (2.35). We say that the entries of $P,Q$ parametric coordinates.
@@
In the following Sections 47, setting some of the parametric coordinates of $\mathfrak{F}_{\mathbb{A}}^{22}$ to constants, or impose linear relations on them, we will obtain several affine varieties whose weighted projectivizations produce examples of prime $\mathbb{Q}$-Fano 3-folds of anti-canonical codimension 4. We refer to Section 8 for another aspect of $\mathfrak{F}_{\mathbb{A}}^{22}$.
@@ 4.1. Definition
We use the following notation for entries of the matrices $P$ and $Q$, by which an group action on $\mathfrak{U}_{\mathbb{A}}^{14}$ will be more visible (see Subsection 4.2):
@@ 4.2. ${\rm SL}_{2}\times{\rm SL}_{2}$-action
4.3. Weights for variables and equations
We assign weights for variables of the polynomial ring $S_{\mathfrak{U}}$ such that all the $9$ equations of $\mathfrak{U}_{\mathbb{A}}^{14}$ are homogeneous. Moreover, we assume that all the variables are not zero allowing some of them are constants. Then it is easy to derive the following relations between the weights of variables of $S_{\mathfrak{U}}$:
@@ 4.3. Weights for variables and equations
In Subsection 4.9, we use the following weights of coordinates:
@@
be the weighted projectivization of $\mathfrak{U}_{\mathbb{A}}^{14}$ by these weights of coordinates. In Subsection 4.9, we show that $\mathfrak{U}_{\mathbb{P}}^{13}$ is a key variety of a prime $\mathbb{Q}$-Fano 3-fold of anticanonical codimension 4 with No.20544. Existence of a positive grading of $\mathfrak{U}_{\mathbb{A}}^{14}$ also plays a role in the proof of Proposition 4.5.
@@ 4.4. Charts and singular locus
For a coordinate *, we call the open subset of $\mathfrak{U}_{\mathbb{A}}^{14}$ with 0$\not$=0∗ ≠ 0 the *-chart. We describe the *-chart such that * is one of the entries of $x,y$, or $s$, $t$.
@@
$({\rm Sing}\,\mathfrak{U}_{\mathbb{A}}^{14})\cap\{y_{1}\not=0\}$$\{y_{1}\not=0,\,x=0,\,s=0,\,M_{y}=O\}$
@@
$({\rm Sing}\,\mathfrak{U}_{\mathbb{A}}^{14})\cap\{y_{2}\not=0\}$$\{y_{2}\not=0,\,x=0,\,s=0,\,M_{y}=O\}$
@@
and it has $c({\rm G}(2,5))$-singularities along $\Delta$, where we call a singularity isomorphic to the vertex of the cone over ${\rm G}(2,5)$ a $c({\rm G}(2,5))$-singularity.
@@
Let $S_{\mathfrak{U}}$ be the polynomial ring over $\mathbb{C}$ whose variables are $s,t$ and the entries of $A,B,C,x,y$. Let $I_{\mathfrak{U}}$ be the ideal of the polynomial ring $S_{\mathfrak{U}}$ generated by the $9$ equations of $\mathfrak{U}_{\mathbb{A}}^{14}$. Set $R_{\mathfrak{U}}:=S_{\mathfrak{U}}/I_{\mathfrak{U}}$. The following assertions hold ::::
@@ 1. $(1)$
    We give nonnegative weights for coordinates of $S_{\mathfrak{U}}$ such that all the equations of $\mathfrak{U}_{\mathbb{A}}^{14}$ are weighted homogeneous, and we denote by $w(*)$ the weight of the monomial *. We denote by $\mathbb{P}$ the corresponding weighted projective space, and by $\mathfrak{U}_{\mathbb{P}}\subset\mathbb{P}$ the weighted projectivization of $\mathfrak{U}_{\mathbb{A}}^{14}$, where we allow some coordinates being nonzero constants (thus $\dim\mathfrak{U}_{\mathbb{P}}$ could be less than $13$). We set
@@ 2. $({1}\text{-}2)$
      The ideal $I_{\mathscr{H}}$ has the following graded minimal $S_{\mathscr{H}}$-free resolution
@@ 4. $(4)$
We may compute the $S_{\mathfrak{U}}$-free resolution (4.7) of $I_{\mathfrak{U}}$ by Singular [DGPS]. For the remaining assertions, the proof of [Tak6, Prop.4.8] works verbatim. ∎
@@ 4.6. Factoriality of $\mathfrak{U}_{\mathbb{A}}^{14}$
By [Tak7, Lem.3.6], it suffices to show that ${\rm Sing}\,(\mathfrak{U}_{\mathbb{A}}^{14}\cap\{x_{1}=0\})$ has codimension $\geq 2$ in $\mathfrak{U}_{\mathbb{A}}^{14}\cap\{x_{1}=0\}$. By the descriptions of charts as in Subsection 4.4, we see that $\mathfrak{U}_{\mathbb{A}}^{14}\cap\{x_{1}=0\}$ is smooth on the $x_{2}$- and the $t$-charts, and is isomorphic to a hypersurface
@@
Using the description of the $x_{1}$-chart as in Subsection 4.4 and Lemma 4.6, the proof of [Tak6, Prop.4.9] work verbatim. ∎
The following corollary can be proved in the same way as the proof of [Tak6, Cor.4.10]:
@@
Let $\mathfrak{U}_{\mathbb{P}}^{13}$ be the weighted projectivization of $\mathfrak{U}_{\mathbb{A}}^{14}$ with some positive weights of coordinates. The following assertions hold:
@@ 4.7. $\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{1}$-fibration
We denote by $\mathbb{A}_{\mathsf{\mathtt{P}}}\simeq\mathbb{A}^{10}$ the affine space whose coordinates are the entries of the matrices $A,B,C$. Note that any equation of $\mathfrak{U}_{\mathbb{A}}^{14}$ is of degree two if we regard the entries of $A,B,C$ as constants. Therefore, considering the variables of the equations of $\mathfrak{U}_{\mathbb{A}}^{14}$ except the entries of $A,B,C$ as projective coordinates, we obtain a $13$-dimensional quasi-projective variety with the same equation as $\mathfrak{U}_{\mathbb{A}}^{14}$. We denote this $13$-dimensional variety by $\widehat{\mathfrak{U}}$, and by $\rho_{\mathfrak{U}}\colon\widehat{\mathfrak{U}}\to\mathbb{A}_{\mathtt{P}}$ the natural projection. Note that the ${\rm SL_{2}}\times{\rm SL}_{2}$-action on $\mathfrak{U}_{\mathbb{A}}^{14}$ defined as in Subsection 4.2 induces those on $\widehat{\mathfrak{U}}$ and those on $\mathbb{A}_{\mathtt{P}}$, and the natural projection $\rho_{\mathfrak{U}}\colon\widehat{\mathfrak{U}}\to\mathbb{A}_{\mathtt{P}}$ is equivariant with respect to these actions. Using these group actions, detail analysis of $\rho_{\mathfrak{U}}\colon\widehat{\mathfrak{U}}\to\mathbb{A}_{\mathtt{P}}$ is possible. We, however, omit details since it is lengthy. Instead, we will describe the restriction of $\rho_{\mathfrak{U}}$ to the subvariety $\mathfrak{S}_{\mathbb{A}}^{8}$ defined in Section 5. Here we only state the following, which can be shown immediately by Corollary 2.31:
@@
Let $\Delta_{\rho_{\mathfrak{U}}}$ be the closed subset of $\mathbb{A}_{\mathtt{P}}$ defined by $D_{\beta}$ as in Remark 2.30. The $\rho_{\mathfrak{U}}$-fiber over a point outside $\Delta_{\rho_{\mathfrak{U}}}$ is isomorphic to ×$\mathbb{P}\times\mathbb{P}^{1}\times\mathbb{P}^{1}.$
@@ 4.8. Terminal singularities
By Proposition 4.4 and the result in Subsection 4.7, we can show the following in the same way as [Tak6, Prop.4.11]:
@@
    The singularities along the $7$-dimensional locus $(\mathsf{S}_{x}\cup\mathsf{S}_{y})\setminus\{x=0,y=0,s=t=0\}$ are $c(G(2,5))$-singularities.
@@
      singularities of $\widetilde{\mathfrak{U}}$ are only $c(G(2,5))$-singularities along the strict transforms of $\mathsf{S}_{x}\cup\mathsf{S}_{y}$, and
@@ 4.9. An example of a $\mathbb{Q}$-Fano threefold
by the weights of coordinates as in Example 4.3. In this subsection, we show the following:
@@
In this proof, we denote $\mathbb{P}(1^{15},2^{2},3)$ by $\mathbb{P}$ for simplicity. Using the equations of $\mathfrak{U}_{\mathbb{P}}^{13}$, we may easily verify that ${\rm Bs}|\mathcal{O}_{\mathbb{P}}(1)|\cap\mathfrak{U}_{\mathbb{P}}^{13}$ consists of the $s$-point and the $t$-point, where the *-point for a coordinate * of $\mathbb{P}$ means the point of $\mathbb{P}$ such that all the coordinates except * are zero. Note that $\dim{\rm Sing\,}\mathfrak{U}_{\mathbb{A}}^{14}$ is less than the codimension of $X$ in $\mathfrak{U}_{\mathbb{P}}^{13}$. Therefore, by the Bertini theorem, we see that $X$ is a smooth $3$-fold outside the $s$-point and the $t$-point. Computing the linear parts of the equations of $X$ at each of the $s$-point and the $t$-point (cf. LPC in [Tak3, Subsec. 5.1]), we see that $X$ has a $1/3(1,1,2)$-singularity at the $s$-point and a $1/2(1,1,1)$-singularity at the $t$-point. By Proposition 4.5 (1-2), we have the following $S_{\mathfrak{U}}$-free minimal resolution of $R_{\mathfrak{U}}$ which is graded with respect to the weights of variables given in Example 4.3:
@@
From this, we see that $-K_{X}=\mathcal{O}_{X}(1)$ and $(-K_{X})^{3}=31/6$. It remains to show that $\rho(X)=1$. By Corollary 4.8, it suffices to check that $X\cap\{x_{1}=0\}$ is a prime divisor. This follows by the Bertini theorem since $\mathfrak{U}_{\mathbb{P}}^{13}\cap\{x_{1}=0\}$ is normal by Proposition 4.5 (3), and the fact that ${\rm Bs}|\mathcal{O}_{\mathbb{P}}(1)|\cap\mathfrak{U}_{\mathbb{P}}^{13}$ consists of the $s$-point and the $t$-point. Hence $X$ is a prime $\mathbb{Q}$-Fano 3-fold of No.20544, and by [Tak6, Proof of Thm.1.2 (1)], $X$ is of anti-canonical codimension $4$. ∎
@@
In a future work, we describe via $\mathfrak{U}_{\mathbb{P}}^{13}$ the Sarkisov link for a prime $\mathbb{Q}$-Fano threefold of No.20544 starting from the weighted blow-up at the unique 1/3(1,1,2)-singularity.

### v3

@@ 2.5. More on FTS and the scheme $\mathfrak{R}$
Explicit descriptions of the quartic form $\widetilde{F}$ and the symplectic form $\omega$ are given as follows (cf.[Cl, (21) in Sec.7, p.118–119],[Fe, p.321], [LM, Prop.5.5]). The proof is same as that given in [Fe, p.321] with the above formulation.
@@
[…]
@@ 2.6. A coordinatization of $\mathfrak{R}$ with parameters
Assume that $N_{x}$ is not identically zero on $V_{x}$. The equations (2.23) as in Theorem 2.18 are reduced to the two equations
@@
[…]
@@
[…]
@@
    There exist the uniquely determined cubic forms $N_{x}^{PQ}(x)$ and $N_{y}^{PQ}(y)$ such that
@@
3. The affine scheme $\mathfrak{F}_{\mathbb{A}}^{22}$
@@
In the affine $26$-space whose coordinates are $s,t,$ and the entries of $x,y,P,Q$, we define $\mathfrak{F}_{\mathbb{A}}^{22}$ to be the scheme with the equations (2.33), (2.34) and (2.35). We call the entries of $P,Q$ the parametric coordinates.
@@
In the following Sections 47, setting some of the parametric coordinates of $\mathfrak{F}_{\mathbb{A}}^{22}$ to constants, or imposing linear relations on them, we will obtain several affine varieties whose weighted projectivizations produce examples of prime $\mathbb{Q}$-Fano 3-folds of anti-canonical codimension 4. We refer to Section 8 for another aspect of $\mathfrak{F}_{\mathbb{A}}^{22}$.
@@ 4.1. Definition
We use the following notation for the entries of the matrices $P$ and $Q$, by which a group action on $\mathfrak{U}_{\mathbb{A}}^{14}$ will be more visible (see Subsection 4.2):
@@ 4.2. ${\rm SL}_{2}\times{\rm SL}_{2}$-action
4.3. Weights on the variables and the equations
We assign weights on the variables of the polynomial ring $S_{\mathfrak{U}}$ such that all the $9$ equations of $\mathfrak{U}_{\mathbb{A}}^{14}$ are homogeneous. Moreover, we assume that all the variables are not zero allowing some of them to be constants. Then it is easy to derive the following relations between the weights on the variables of $S_{\mathfrak{U}}$:
@@ 4.3. Weights on the variables and the equations
In Subsection 4.9, we use the following weights on the coordinates:
@@
be the weighted projectivization of $\mathfrak{U}_{\mathbb{A}}^{14}$ by these weights on the coordinates. In Subsection 4.9, we show that $\mathfrak{U}_{\mathbb{P}}^{13}$ is a key variety of a prime $\mathbb{Q}$-Fano 3-fold of anticanonical codimension 4 with No.20544. Existence of a positive grading of $\mathfrak{U}_{\mathbb{A}}^{14}$ also plays a role in the proof of Proposition 4.5.
@@ 4.4. Charts and singular locus
For a coordinate *, we call the open subset of $\mathfrak{U}_{\mathbb{A}}^{14}$ with $\not=0$ the *-chart. We describe the *-chart such that * is one of the entries of $x,y$, or $s$, $t$.
@@
$({\rm Sing}\,\mathfrak{U}_{\mathbb{A}}^{14})\cap\{y_{1}\not=0\}$ $=\{y_{1}\not=0,\,x=0,\,s=0,\,M_{y}=O\}$
@@
$({\rm Sing}\,\mathfrak{U}_{\mathbb{A}}^{14})\cap\{y_{2}\not=0\}$ $=\{y_{2}\not=0,\,x=0,\,s=0,\,M_{y}=O\}$
@@
and it has $c({\rm G}(2,5))$-singularities along $\Delta$, where we call a singularity isomorphic to the vertex of the cone over ${\rm G}(2,5)$ a $c({\rm G}(2,5))$-singularity (it is easy to verify that this is a terminal singularity by computing the discrepancy of the exceptional divisor of the simple blowing up at the point).
@@
Let $S_{\mathfrak{U}}$ be the polynomial ring over $\mathbb{C}$ whose variables are $s,t$ and the entries of $A,B,C,x,y$. Let $I_{\mathfrak{U}}$ be the ideal of the polynomial ring $S_{\mathfrak{U}}$ generated by the $9$ equations of $\mathfrak{U}_{\mathbb{A}}^{14}$. Set $R_{\mathfrak{U}}:=S_{\mathfrak{U}}/I_{\mathfrak{U}}$. The following assertions hold ::
@@ 1. $(1)$
    We give nonnegative weights for the coordinates of $S_{\mathfrak{U}}$ such that all the equations of $\mathfrak{U}_{\mathbb{A}}^{14}$ are weighted homogeneous, and we denote by w()w(*) the weight on the monomial *. We denote by $\mathbb{P}$ the corresponding weighted projective space, and by $\mathfrak{U}_{\mathbb{P}}\subset\mathbb{P}$ the weighted projectivization of $\mathfrak{U}_{\mathbb{A}}^{14}$, where we allow some coordinates being nonzero constants (thus $\dim\mathfrak{U}_{\mathbb{P}}$ could be less than $13$). We set
@@ 2. $({1}\text{-}2)$
      The ideal $I_{\mathfrak{U}}$ has the following graded minimal $S_{\mathfrak{U}}$-free resolution
@@ 4. $(4)$
We may compute the $S_{\mathfrak{U}}$-free resolution (4.1) of $I_{\mathfrak{U}}$ by Singular [DGPS]. For the remaining assertions, the proof of [Tak4, Prop.4.8] works verbatim. ∎
@@ 4.6. Factoriality of $\mathfrak{U}_{\mathbb{A}}^{14}$
By [Tak5, Lem.3.6], it suffices to show that ${\rm Sing}\,(\mathfrak{U}_{\mathbb{A}}^{14}\cap\{x_{1}=0\})$ has codimension $\geq 2$ in $\mathfrak{U}_{\mathbb{A}}^{14}\cap\{x_{1}=0\}$. By the descriptions of charts as in Subsection 4.4, we see that $\mathfrak{U}_{\mathbb{A}}^{14}\cap\{x_{1}=0\}$ is smooth on the $x_{2}$- and the $t$-charts, and is isomorphic to a hypersurface
@@
Using the description of the $x_{1}$-chart as in Subsection 4.4 and Lemma 4.6, the proof of [Tak4, Prop.4.9] works verbatim. ∎
The following corollary can be proved in the same way as the proof of [Tak4, Cor.4.10]:
@@
Let $\mathfrak{U}_{\mathbb{P}}^{13}$ be the weighted projectivization of $\mathfrak{U}_{\mathbb{A}}^{14}$ with some positive weights on the coordinates. The following assertions hold:
@@ 4.7. $\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{1}$-fibration
We denote by $\mathbb{A}_{\mathsf{\mathtt{P}}}\simeq\mathbb{A}^{10}$ the affine space whose coordinates are the entries of the matrices $A,B,C$. Note that any equation of $\mathfrak{U}_{\mathbb{A}}^{14}$ is of degree two if we regard the entries of $A,B,C$ as constants. Therefore, considering the variables of the equations of $\mathfrak{U}_{\mathbb{A}}^{14}$ except the entries of $A,B,C$ as projective coordinates, we obtain a $13$-dimensional quasi-projective variety with the same equation as $\mathfrak{U}_{\mathbb{A}}^{14}$. We denote this $13$-dimensional variety by $\widehat{\mathfrak{U}}$, and by $\rho_{\mathfrak{U}}\colon\widehat{\mathfrak{U}}\to\mathbb{A}_{\mathtt{P}}$ the natural projection. Note that the ${\rm SL_{2}}\times{\rm SL}_{2}$-action on $\mathfrak{U}_{\mathbb{A}}^{14}$ defined as in Subsection 4.2 induces those on $\widehat{\mathfrak{U}}$ and those on $\mathbb{A}_{\mathtt{P}}$, and the natural projection $\rho_{\mathfrak{U}}\colon\widehat{\mathfrak{U}}\to\mathbb{A}_{\mathtt{P}}$ is equivariant with respect to these actions. Using these group actions, detail analysis of $\rho_{\mathfrak{U}}\colon\widehat{\mathfrak{U}}\to\mathbb{A}_{\mathtt{P}}$ is possible. We, however, omit the details since they are lengthy. Instead, we will describe the restriction of $\rho_{\mathfrak{U}}$ to the subvariety $\mathfrak{S}_{\mathbb{A}}^{8}$ defined in Section 5. Here we only state the following, which can be shown immediately by Corollary 2.31:
@@
Let $\Delta_{\rho_{\mathfrak{U}}}$ be the closed subset of $\mathbb{A}_{\mathtt{P}}$ defined by $D_{\beta}$ as in Remark 2.30. The $\rho_{\mathfrak{U}}$-fiber over a point outside $\Delta_{\rho_{\mathfrak{U}}}$ is isomorphic to $\mathbb{P}\times\mathbb{P}^{1}\times\mathbb{P}^{1}.$
@@ 4.8. Terminal singularities
By Proposition 4.4 and the result in Subsection 4.7, we can show the following in the same way as [Tak4, Prop.4.11]:
@@
    The singularities along the $7$-dimensional locus $(\mathsf{S}_{x}\cup\mathsf{S}_{y})\setminus\{x=0,y=0,s=t=0\}$ are $c({\rm G}(2,5))$-singularities (cf. Proposition 4.4).
@@
      singularities of $\widetilde{\mathfrak{U}}$ are only $c({\rm G}(2,5))$-singularities along the strict transforms of $\mathsf{S}_{x}\cup\mathsf{S}_{y}$, and
@@ 4.9. An example of a $\mathbb{Q}$-Fano threefold
by the weights on the coordinates as in Example 4.3. In this subsection, we show the following:
@@
In this proof, we denote $\mathbb{P}(1^{15},2^{2},3)$ by $\mathbb{P}$ for simplicity. Using the equations of $\mathfrak{U}_{\mathbb{P}}^{13}$, we may easily verify that ${\rm Bs}|\mathcal{O}_{\mathbb{P}}(1)|\cap\mathfrak{U}_{\mathbb{P}}^{13}$ consists of the $s$-point and the $t$-point, where the *-point for a coordinate * of $\mathbb{P}$ means the point of $\mathbb{P}$ such that all the coordinates except * are zero. Note that $\dim{\rm Sing\,}\mathfrak{U}_{\mathbb{A}}^{14}$ is less than the codimension of $X$ in $\mathfrak{U}_{\mathbb{P}}^{13}$. Therefore, by the Bertini theorem, we see that $X$ is a smooth $3$-fold outside the $s$-point and the $t$-point. Computing the linear parts of the equations of $X$ at each of the $s$-point and the $t$-point (cf. LPC in [Tak7, Subsec.2.8.1]), we see that $X$ has a $1/3(1,1,2)$-singularity at the $s$-point and a $1/2(1,1,1)$-singularity at the $t$-point. By Proposition 4.5 (1-2), we have the following $S_{\mathfrak{U}}$-free minimal resolution of $R_{\mathfrak{U}}$ which is graded with respect to the weights on the variables given in Example 4.3:
@@
From this, we see that $K_{X}=\mathcal{O}_{X}(1)$ and $(-K_{X})^{3}=31/6$. It remains to show that $\rho(X)=1$. By Corollary 4.8, it suffices to check that $X\cap\{x_{1}=0\}$ is a prime divisor. This follows by the Bertini theorem since $\mathfrak{U}_{\mathbb{P}}^{13}\cap\{x_{1}=0\}$ is normal by Lemma 4.6, and the fact that ${\rm Bs}|\mathcal{O}_{\mathbb{P}}(1)|\cap\mathfrak{U}_{\mathbb{P}}^{13}$ consists of the $s$-point and the $t$-point. Hence $X$ is a prime $\mathbb{Q}$-Fano 3-fold of No.20544, and by [Tak7, Subsec.2.5, Proof of Thm.1 (1)], $X$ is of anti-canonical codimension $4$. ∎
@@
In a future work, we will describe via $\mathfrak{U}_{\mathbb{P}}^{13}$ the Sarkisov link for a prime $\mathbb{Q}$-Fano threefold of No.20544 starting from the weighted blow-up at the unique 1/3(1,1,2)-singularity.